OpenStudy (anonymous):

The math team sold 42 hamburger meals and 15 ribs meals at the pep rally for a total of $243. At the game that night they sold 36 hamburger meals and 30 ribs meals for a total of $294. What is the cost of 1 rib meal?

OpenStudy (mertsj):

42h+15r=243 36h+30r=294

OpenStudy (anonymous):

Let one hamburger meal costs x and one rib meal costs y.. So, according to the question : 42x + 15y = 243 -----------1 36x + 30y = 294 The lower equation can be more simplified by dividing by 2 both sides: 18x + 15y = 147 ------------2 Therefore, subtract equation 1 and equation 2, 24x = 96 So, x = 4 Put value of x in any one equation to get the value of y: 15y = 147 - 18x 15y = 147 - 18(4) 15y = 147 - 72 15y = 75 Therefore y = 5.. Therefore, cost of 1 rib meal (y) is $5...